1.    Balance each of the following equations according to the half-reaction method:

a) Br2(l) + SO2(g) ⟶ Br(aq) + SO42−(aq)                               (in acid)

b) Cl2(g) + OH−(aq) ⟶ Cl(aq) + ClO3(aq)                            (in base)

 



1
Expert's answer
2022-01-04T01:25:19-0500

a) Br2(l) + 2e- =  2Br(aq

 SO2(g) + 2H2O (l) =  SO42−(aq) + 4H+ (aq) +2e-


Total: Br2(l) + SO2(g) + 2H2O (l) =  2Br(aq) + SO42−(aq) + 4H+ (aq)


b) Cl2(g) + 2e- =  2Cl(aq

 Cl2(g) + 12OH(aq) = 2ClO3(aq) + 6H2O (l) + 10e-


Total:  3Cl2(g) + 6OH(aq) = 5Cl(aq) + ClO3(aq) + 3H2O (l)